the sum of two intergers is -33 and their product is 162.develop a quadratic equation and then solve to find the value of the two intergers?
Found 3 solutions by ikleyn, addingup, Alan3354:
Answer by ikleyn(52780)
Answer by addingup(3677)

x + y = -33
y = (-33 - x)
therefore:
(-33 - x) + x = -33
and:
(-33 - x) * x = 162
Now let's set up the quadratic and then solve:
-33x - x^2 = 162
-33x - x^2 -162 = 0
let's write it in the standard form ax^2 + bx + c = 0:
-x^2 - 33x - 162 = 0
(−x − 6)(x + 27) = 0
−x − 6 = 0 or x + 27 = 0
x = −6 or x = −27
Our two numbers are -6 and -27

Answer by Alan3354(69443)

integers.
-----
there are no intergers.